# Conceptualization, Methodology, Software, Writing - Original draft

### v1

@@ 2.1 Knowledge Graph Construction from Text
LLMs enable zero-shot and few-shot extraction without task-specific fine-tuning [zhu2024llms, carta2023iterative]. However, LLM-based extraction introduces challenges: hallucination of factually incorrect triples, lack of structural consistency mechanisms, and bias toward frequent relations at the expense of rare but important types [ji2023hallucination, agrawal2024knowledge]. These limitations motivate complementary validation mechanisms to verify LLM-generated extractions.
@@ 3.2.3 Hierarchical Relation Schema
Table 1 presents the complete hierarchical relation schema, organizing all 89 fine-grained relation types into 8 coarse-grained categories. This comprehensive categorization supports the model in distinguishing subtle semantic differences between relations such as locatedIn and distributedIn.
Table 1: Hierarchical relation schema with 8 coarse-grained categories and 89 fine-grained relation types. Relations marked with $\dagger$ are long-tail types (frequency <5<5).
Category Fine-grained Relations Count
Hierarchy & Composition belongsTo, contains, composedOf, partOf 4
Total 89
@@ 3.3.1 Single-Pass Hierarchical Extraction
[…]
@@
[…]
@@ 3.3.2 Schema-Constrained Prompt Design
The prompt embeds the complete hierarchical schema (detailed in Table 1), including: entity type definitions with examples, the 8 coarse-grained categories with semantic descriptions, and fine-grained relations organized under each category.
@@ 3.3.3 Out-of-Schema Relation Remapping
[…]
@@ 4.1.2 Baselines
[…]
@@ 4.1.4 Implementation Details
Table 5: Implementation details and hyperparameter configurations.
@@
     Strategy Hierarchical (8 coarse $\to$ 89 fine)
     Chunk Size 2,000 chars (200 overlap)
Semantic Initialization
     Encoder Chinese-RoBERTa-wwm-ext (frozen)
     Projection Linear ($768\to 1024$)
@@
     Embedding Dim 512 (Complex Space)
     Negative Sampling Self-adversarial ($\alpha=1.0$)
     Update Strategy Warm-start each iteration
@@
     Retry Limit $K=3$
Channel 2: Triple Integration
     High Confidence Top 30% (Accept)
     Medium Confidence Middle 45% (Verify)
     Low Confidence Bottom 25% (Reject)
@@ 4.4 Convergence Analysis (RQ4)
Figure 6 tracks performance and knowledge graph scale across iterations.
Figure 6: Convergence over iterations: Micro-F1 (%) and cumulative valid triples.
T=1 (Initialization). The framework establishes the initial knowledge graph with 21.5% Micro-F1 and 341 validated triples. At this stage, the RotatE model is undertrained, leading to conservative validation that filters some valid triples.
T=2 (Foundation Building). Channel 1’s Evidence-Guided CoT Feedback begins recovering erroneously rejected triples. Micro-F1 improves to 27.2% (+5.7 points), with the knowledge graph expanding to 436 triples.
T=3 (Accelerated Growth). The bidirectional collaboration enters its productive phase. Channel 2’s Active Selection identifies high-value samples that expand coverage, while Channel 1 continues refining extraction quality. Micro-F1 reaches 32.1% (+4.9 points), with 572 validated triples.
T=4 (Maturation). The framework achieves peak performance with 36.79% Micro-F1 (+4.69 points) and 727 validated triples. The continued improvement reflects the cumulative effect of progressively refined models.
[…]
@@ 5.2 Limitations
[…]
However, knowledge graph construction is inherently an offline process (i.e., performed as a one-time preprocessing step before deployment) rather than a real-time query service. Therefore, this latency is architecturally acceptable. The significant gains in precision and the reduction of hallucinated relations justify the additional computational investment during the construction phase.

### v2

# arXiv:2602.02090v2[cs.CL] 27 Feb 2026

@@ 2.1 Knowledge Graph Construction from Text
LLMs enable zero-shot and few-shot extraction without task-specific fine-tuning [zhu2024llms, carta2023iterative]. However, LLM-based extraction introduces challenges: hallucination of factually incorrect triples, lack of structural consistency mechanisms, and bias toward frequent relations at the expense of rare but important types [ji2023survey, agrawal2024knowledge]. These limitations motivate complementary validation mechanisms to verify LLM-generated extractions.
@@ 3.2.2 Entity Type Hierarchy
Table 1: Coarse-grained relation categories with representative fine-grained types.
Category Representative Relations #
Definition & Naming definedAs, fullNameOf, aliasOf 4
Hierarchy & Composition belongsTo, contains, partOf 4
Spatiotemporal locatedIn, timeRangeOf, covers 11
Quantitative hasValue, hasUnit, meanValueOf 14
Trend & Change trendOf, changeRateOf, growthRateOf 6
Provenance & Method dataSourceOf, usesMethod, citedFrom 17
Causality & Impact causes, affects, promotes, mitigates 10
Application & Status usedFor, statusOf, contributesToSDG 23
@@ 3.2.3 Hierarchical Relation Schema
Table 2 summarizes the 8 categories with representative relations. The complete schema with all 89 fine-grained relations is provided in Appendix A.
Table 2: Coarse-grained relation categories with representative fine-grained types.
Category Representative Relations #
Definition & Naming definedAs, fullNameOf, aliasOf 4
Hierarchy & Composition belongsTo, contains, partOf 4
Spatiotemporal locatedIn, timeRangeOf, covers 11
Quantitative hasValue, hasUnit, meanValueOf 14
Trend & Change trendOf, changeRateOf, growthRateOf 6
Provenance & Method dataSourceOf, usesMethod, citedFrom 17
Causality & Impact causes, affects, promotes, mitigates 10
Application & Status usedFor, statusOf, contributesToSDG 23
@@ 3.3.1 Single-Pass Hierarchical Extraction
8:  $\mathcal{P}\leftarrow\text{BuildHierarchicalPrompt}(c,\mathcal{O|}|)$ $\triangleright$ Coarse-to-fine
9:  $\mathcal{T}_{c}\leftarrow\mathcal{M}_{\text{LLM|}|}(\mathcal{P})$ $\triangleright$ Output: $(h,r,t,e,c)$
@@
19:   $\mathcal{T}_{\text{oos}}\leftarrow\mathcal{T}_{\text{oos}}\cup\{\tau\}$ $\triangleright$ Out-of-schema
@@ 3.3.2 Schema-Constrained Prompt Design
The prompt embeds the complete hierarchical schema (Appendix A), including: entity type definitions with examples, the 8 coarse-grained categories with semantic descriptions, and fine-grained relations organized under each category.
@@ 3.3.3 Out-of-Schema Relation Remapping
If remapping fails, the tuple is discarded. This conservative strategy prioritizes precision over recall, as structurally invalid triples would degrade downstream KGE training.
@@ 4.1.2 Baselines
We compare our framework against the following baselines:
  • LLM One-shot: DeepSeek-V3 with a single demonstration example.
  • LLM Few-shot: DeepSeek-V3 with 5-shot in-context examples covering diverse relation types.
  • OpenIE + Semantic Mapping: Open information extraction followed by semantic clustering to map extracted relations to the 87 predefined types.
  • Schema-constrained (No Iter.): Our extraction pipeline excluding the bidirectional refinement module, serving as a single-pass baseline.
@@ 4.1.4 Implementation Details
Table 6: Implementation details.
@@
     Strategy Hierarchical (8-class → 87-class)
RoBERTa (Frozen)
     Model Chinese-RoBERTa-wwm-ext
     Training Data 2023 + 2024 entities (for alignment)
     Usage Semantic Init + MC-Dropout
@@
     Embedding Dim 512
     Initial Training 2023 triples (1,548)
     Update Warm-start each iteration
@@
Channel 2: Active Selection
     MC-Dropout Runs 5
     Selection High structure + High uncertainty
@@ 4.4 Convergence Analysis (RQ4)
Previous experiments evaluated final output quality via F1 scores. This section analyzes the behavioral characteristics of the bidirectional collaboration mechanism across iterations. To enable consistent evaluation across all nine rounds, we adopt precision and validated triple count as convergence metrics, as these can be computed at each iteration without requiring full re-evaluation on the test set. We extend the iteration rounds to 9 to comprehensively observe convergence behavior and potential overfitting. Table 10 and Figure 6 present the detailed metrics and convergence curves.
Table 10: Validation metrics across iteration rounds.
Round #Validated Triples Precision (%)
1 1100 29.1
2 1300 31.5
3 1530 34.0
4 1965 34.9
5 2060 36.7
6 1905 35.7
7 1700 33.5
8 1518 30.6
9 1400 28.0
The iterative process exhibits three distinct phases:
Figure 6: Convergence over iterations: validated triple count and precision (%).
Collaborative Growth (Round 1–4). During early iterations, the bidirectional collaboration mechanism progressively takes effect. The RotatE validator strengthens its structural scoring capability with increasing training data, while the CoT evidence feedback recovers correctly extracted triples that were conservatively rejected in earlier rounds. The synergy drives validated triples from 1,100 to 1,965 and precision from 29.1% to 34.9%, demonstrating simultaneous improvement in both quantity and quality. Notably, the Round 4 precision (34.9%) closely aligns with the overall framework precision (34.84%) reported in Section 4.2, confirming the consistency of our evaluation.
Peak Performance (Round 5–6). The framework reaches optimal performance at Round 5, with 2,060 validated triples and 36.7% precision. At this stage, RotatE and CoT feedback achieve their best collaboration: RotatE accurately identifies structurally plausible triples while CoT provides semantic-level verification for borderline cases. Round 6 shows a slight decline (1,905 triples, 35.7% precision) but remains at a high level, indicating a brief plateau before degradation.
Overfitting Decay (Round 7–9). From Round 7 onward, performance degrades continuously. Validated triples decrease from 1,700 to 1,400, and precision drops from 33.5% to 28.0%. We hypothesize this degradation is primarily driven by RotatE overfitting: after excessive training iterations, the model becomes overly sensitive to specific patterns in historical data, leading to distorted scoring criteria that both reject valid borderline triples and memorize noise patterns. Notably, Round 9 precision (28.0%) falls below Round 1 (29.1%), indicating that over-iteration not only fails to yield gains but erases earlier improvements.
Validation of Early Stopping. These results validate the necessity of early stopping. While Round 5 achieves peak precision (36.7%), the declining trend from Round 6 suggests that stopping at Round 4 ($T=4$) offers better stability margins. The Round 4 configuration balances two advantages: its precision (34.9%) is already close to the peak with limited loss, and it provides a sufficient safety buffer against overfitting. The final performance reported in Section 4.2 (Micro-F1 = 36.79%) is based on $T=4$, which achieves a favorable trade-off between extraction quality and system robustness.
Key Insights. (1) Bidirectional collaboration achieves simultaneous quantity-quality improvement in early iterations, validating the synergy between structural validation and evidence-guided feedback. (2) A clear performance peak exists (Round 5), beyond which overfitting causes continuous degradation. (3) Over-iteration can regress performance below initial levels, making early stopping ($T=4$) essential for stable output.
@@ 5.2 Limitations
Computational Overhead.
The iterative refinement process requires multiple rounds of LLM inference and embedding model updates. For resource-constrained settings, a trade-off between iteration depth and computational cost must be considered.
@@ 6 Conclusion
Appendix A Complete Hierarchical Relation Schema
Table 11 presents the complete hierarchical relation schema with all 89 fine-grained relation types. Relations marked with $\dagger$ are long-tail types with frequency <5<5 in our annotated dataset.
Table 11: Complete hierarchical relation schema. Relations marked with $\dagger$ are long-tail types (frequency <5<5).
Category Fine-grained Relations Count
Hierarchy & Composition belongsTo, contains, composedOf, partOf 4
Total 89